OpenStudy (anonymous):

two planes are 7700 miles apart, and their speeds differ by 300 mph. They travel toward each other and meet in 7 hours. Find the speed of the slower plane.

OpenStudy (amistre64):

hmm ..... 7700 - t -(t+300) = 0, perhaps? its worth a shot

OpenStudy (amistre64):

7700 - t -t -300 = 0 7400 - 2t = 0 7400 = 2t 7400/2 = t = 3700 miles .... but thats prolly a side effect of what I was thinking to begin with since i measured distance instead of time

OpenStudy (anonymous):

Which one should I do then? :)

OpenStudy (amistre64):



OpenStudy (anonymous):

Okay hold on.. I'm going to try it out.

OpenStudy (amistre64):

we equate the speeds and cover the distance 7700 = 7s + 7(s+300) 7700 = 7s + 7s+2100 7700 = 14s +2100 -2100 -2100 ------------------- 5600 = 14s s = 5600/14, but id dbl chk it forst

OpenStudy (amistre64):

if thats done right; the one plane is going 400mph while the other is going 700mph the distance covered when added should be 7700 miles

OpenStudy (anonymous):

I got 400?

OpenStudy (amistre64):

7700 = 7(400) + 7(700) 7700 = 2800 + 4900 7700 = 7700 so yes, the slower is going 400

OpenStudy (anonymous):

Yes! I got it! Your the best :)) Thankyou!

OpenStudy (amistre64):

yep, youre welcome; and i tend to think of myself as the penultimate mathmatiker lol

OpenStudy (anonymous):

I believe it! I would have never gotten that without you. Whats your name?

OpenStudy (amistre64):

.... amistre64

OpenStudy (anonymous):

No like you REAL name. Like mine is Joyce :) lol